OpenStudy (anonymous):

What percent of 10 is 6?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

6/10% = 0.6% 6/10 = 3/5 = 60% a difference of 4%

OpenStudy (anonymous):

if you really care about the calculations, its 6/10 *100=60

OpenStudy (anonymous):

Woe !! ha

OpenStudy (anonymous):

Thanks